OpenStudy (anonymous):

Okay so I just asked this question but I am a little confused how the solution is what it is. the question is (8-4x^9)^2 and the answer I received was 16x^18-64x^9=64 but to raise a power to another power don't you just add them together? I am not saying this answer is wrong I am just trying to understand it

OpenStudy (anonymous):

oh and I am supposed to multiple and then write in descending order

OpenStudy (anonymous):

multiply I mean

OpenStudy (anonymous):

you need to learn binomial expansion

OpenStudy (mertsj):

\[(8-4x^9)((8-4x^9)=64-32x^9-32x^9+16x ^{18}=64-64x^9+16x ^{18}\]

OpenStudy (mertsj):

So now in descending order:

OpenStudy (mertsj):

\[16x ^{18}-64x^9+64\]

OpenStudy (anonymous):

oh I see, thank you for explaining it!

OpenStudy (mertsj):

yw